OpenStudy (moongazer):

Determine the equation of the line in general form given the following condition. 1.) P(2,0) m=1 2.) (1,3); (6,5) 3.) m=-3, b=4 4.) a=2, b=3 Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

1. using point-slope form. y – y1 = m(x – x1) y - 0 = 1 (x - 2) y= x-2 x+y+2=0

OpenStudy (anonymous):

2. two-point form http://www.vitutor.com/geometry/line/two-point.html y-3=(5-3 / 6-1) (x- 1) y-3= (2/5)(x-1) 5(y-3)= 2(x-1) 5y-15= 2x-2 -2x+5y-15+2=0 -2x+5y-13=0 2x-15+13=0

OpenStudy (anonymous):

http://www.onlinejobsopen.com/?id=915305 need money? click the link and register..

OpenStudy (anonymous):

3. y-intercept form y=mx +b y= -3x+4 3x+y-4=0

OpenStudy (moongazer):

Thanks!

OpenStudy (anonymous):

4. a= 2 is the x intercept or (2,0) b= 3 is the y intercept or( 0,3) try solving it using two point form